### v1

# ToolHaystack: Stress-Testing Tool-Augmented Language Models in Realistic Long-Term Interactions

Abstract
Large language models (LLMs) have demonstrated strong capabilities in using external tools to address user inquiries. However, most existing evaluations assume tool use in short contexts, offering limited insight into model behavior during realistic long-term interactions. To fill this gap, we introduce ToolHaystack, a benchmark for testing the tool use capabilities in long-term interactions. Each test instance in ToolHaystack includes multiple tasks execution contexts and realistic noise within a continuous conversation, enabling assessment of how well models maintain context and handle various disruptions. By applying this benchmark to 14 state-of-the-art LLMs, we find that while current models perform well in standard multi-turn settings, they often significantly struggle in ToolHaystack, highlighting critical gaps in their long-term robustness not revealed by previous tool benchmarks. [^1] Our code and data are available at https://github.com/bwookwak/ToolHaystack
Corresponding author.
ToolHaystack: Stress-Testing Tool-Augmented Language Models
in Realistic Long-Term Interactions
Beong-woo Kwak1  Minju Kim1  Dongha Lim1  Hyungjoo Chae1 Dongjin Kang1Sunghwan Kim1Dongil Yang1Jinyoung Yeo1 1Yonsei University {bwoo.kwak,jinyeo}@yonsei.ac.kr
1 Introduction
Recent breakthroughs in Large Language Models (LLMs) have transformed their role toward LLM agents; they can execute real-world tasks such as managing financial transactions and scheduling appointments (Yao et al., 2022; Shinn et al., 2023). Tool-augmented language model (TALM) benchmarks (Qin et al., 2023; Li et al., 2023; Chen et al., 2024) have played a critical role in evaluating tool-use capabilities based on simplified instruction-following setups where they autonomously execute external tools to address user goals.
Figure 1: ToolHaystack addresses long-term interactions that include evolving goals, semantic noise, and fragmented context.
Benchmarks Multi-turn Long Interaction Granular Complexity Distractor Tasks Avg. Turns ToolBench X X X X 1 Fail-TaLMs X X X X 1 RoTBench X X X 1 ComplexFuncBench X X X X 1 ToolDial X X X 4.5 $\tau$-Bench X X X MMTB X X 5.3 HammerBench X X 2.2 BFCL-v3 X X 4.2 ToolHaystack 32
Table 1: Comparison of TALM benchmarks across key evaluation dimensions. Long Interaction captures extended conversations where long-term dependency plays a critical role; Granular Complexity denotes supporting fine-grained evaluation with varying difficulty levels; Distractor Tasks measure robustness under interleaved irrelevant tasks. The average number of turns $\tau$-Bench is marked as ‘–’ as it does not use static dialogues for evaluation.
Despite these advancements, the scope of most existing work is still limited short interaction, which fail to capture the behavior of LLM agents in long-term interactions. The queries of most existing benchmarks are simple (e.g.“Book a restaurant named ….”) which can be resolved within one or two exchanges. In contrast, many real-world tasks, such as travel planning or project coordination, cannot be resolved in short interactions (Wu et al., 2025; Maharana et al., 2024) or steps (Hayati et al., 2025; Press et al., 2023). For example, the user may want to perform tasks based on previous requests without restating all the details (e.g.“Cancel the reservation you made for my birthday.”) or request to collaborate with the agents across long time frame (e.g.“Help me writing a research paper.”). These tasks inherently require interactions that unfold and might evolve over time.
While recent efforts have extended test scenarios through multi-turn interactions (Li et al., 2023) and API response simulations (Yan et al., 2024; Zhong et al., 2025), these approaches fall short in capturing the full complexity of long-term interactions. In real-world contexts, continuous exchanges between the user, the agent, and the external environment introduce various contextual noises—such as interleaved task flows (Castillo-Bolado et al., 2024), redundant information (Bai et al., 2024), and shifting user goals (Yu et al., 2025)—that accumulate in the agent’s context (Fig.1). The current lack of benchmarks that rigorously evaluate TALMs’ robustness under such conditions limits their applicability to real-world scenarios, especially for long-term use cases like personal assistants.
In this paper, we aim to evaluate TALMs in realistic long-term interaction scenarios. To this end, we introduce ToolHaystack, a realistic and composable test suite for TALMs. Specifically, we define long-term agent context as a highly noisy context where multiple user goals and partial task execution histories are entangled in a single context. Test instances are structured to simulate natural task flows by interleaving task-relevant ("needles") and irrelevant contexts ("haystack") [^2] We draw the analogy from the classic needle in a haystack problem (Kamradt, 2023) to emphasize the noisy nature of long-term interaction., posing unique challenges of identifying key information among distractions accumulated over time. Based on the structure, ToolHaystack simulates six realistic scenarios derived from three core challenges (context recall, information shift, and missing context), each with two difficulty levels.
We systematically evaluate 14 highly capable LLMs with tool-use capabilities, covering both open-source and proprietary models on ToolHaystack. Our findings reveal that even highly capable LLMs with advanced long-context modeling (Bai et al., 2024; An et al., 2024) still struggle to maintain consistency and robustness under long-term scenarios—suggesting that contemporary benchmarks may overestimate their readiness and reliability for real-world deployment. Beyond general performance, we also conduct detailed ablation studies to uncover what factors influence long-term success and where current models fail.
Our main contributions are as follows:
  • We are the first to explicitly formalize and highlight the critical challenge of long-term interaction robustness for TALMs.
  • We introduce ToolHaystack, a benchmark to stress-test LLM agents in long-term scenarios.
  • We systematically evaluate 14 state-of-the-art LLM agents, showing key failure modes in long-term interactions and guidance for their robust real-world deployment.
2 Related Work
Realistic Evaluation of TALMs.
To better approximate real-world conditions, benchmarks such as ToolDial (Shim et al., 2025), HammerBench (Wang et al., 2024), and $\tau$-Bench (Yao et al., 2025) introduce more naturalistic, multi-turn dialogues. Recent efforts like RoTBench (Ye et al., 2023) and FAIL-TALM (Treviño et al., 2025) examine robustness under adversarial or erroneous conditions. However, these benchmarks remain largely constrained to single-session or short-horizon tasks, offering limited insight into TALM behavior over extended interactions.
Long-context Evaluation of TALMs.
Recent benchmarks such as MMTB (Yu et al., 2025), BFCL (Yan et al., 2024), and ComplexFuncBench (Zhong et al., 2025) have extended TALM evaluation into longer contexts and compositional tool use. However, they often rely on clean task structures with highly coherent subtasks (< 5) or simulate long context by concatenating tool outputs.
In contrast, ToolHaystack provides a long-context TALM benchmark with interleaved task flows and contextual noise. This exposes novel failure modes that are underexplored in prior evaluations. Tab.1 compares ToolHaystack with existing TALM benchmarks. Further discussions on related work are presented in Appendix A.
3 Rethinking Evaluation of TALMs
Figure 2: Illustration of the scenarios in ToolHaystack (CR, IS, MC).
Most LLM benchmarks are rapidly saturated by frontier models (Sirdeshmukh et al., 2025; Phan et al., 2025). Likewise, previous TALM evaluation frameworks might have limited lifespans. This underscores a constant need for more realistic, challenging and adaptive evaluation frameworks that can keep pace with model advancements. We highlight that long-term interactions go beyond simple extension of standard multi-turn evaluation.
3.1 Long-Term & Multi-Turn Tool-use
The key distinction lies not just in the number of turns, it reflects the inherent complexity of real-world tasks (Fig.1):
  • Fragmented context: Relevant information is scattered across histories. The context often involves the execution history of multiple tasks.
  • Evolving goals: User goals may shift over time. Models should adapt their response accordingly.
  • Semantic noise accumulation: Irrelevant but plausible content such as tool responses from unrelated tasks can mislead models.
3.2 Task Definition
Let $\mathcal{T}=\{\text{tool}_{1},\dots,\text{tool}_{k}\}$ be a set of tools, each taking a structured query as input and returning a result. A task session consists of a sequence of interactions between the TALM, user, and environment. Let $S=\{s_{1},\dots,s_{T}\}$ be the set of sessions, where each $s_{t}=\{x_{t}^{1},a_{t}^{1},r_{t}^{1},\dots,x_{t}^{n},a_{t}^{n},r_{t}^{n}\}$ comprises user inputs $x_{t}^{i}$, model outputs $a_{t}^{i}$ (natural language response or tool calls), and tool response $r_{t}^{i}$ returned by the tool call (if applicable). Given the session history $H_{t}=\{s_{1},\dots,s_{t-1}\}$ and the current session, the model chooses an action $a_{t}^{i}$ from the action space where a tool action is represented as a pair (tool_name, arguments).
4 ToolHaystack: A Realistic and Composable Test Suite for Long-Term Tool Utilization
We introduce ToolHaystack, a benchmark to stress-test TALM’s abilities in long-horizon tasks.
4.1 Test Structure
The key principles of ToolHaystack are: (1) The interaction should be aligned with how real-world users communicate with agents; and (2) The difficulty and contextual interference levels should be adjustable to keep the pace of model improvements. However, annotating such test is difficult and time-consuming even for human experts.
To address this issue, we design a novel test structure for ToolHaystack, inspired by Needle-In-A-Haystack (NIAH) dataset (Kamradt, 2023). The key idea is to interleave different task sessions as a context instead of isolating task sessions (Fig.1). In ToolHaystack, needles are defined as parts of the context that are essential to address the target task. The term haystack refers to the remaining (usually irrelevant) text, that acts as a distractor. This modularized structure enables us to control the difficulty of test instances by manipulating how we compose this haystack. For example, increasing the number of task sessions in the haystack can naturally increase the task difficulty.
4.2 Test Scenarios
To rigorously evaluate the capabilities of TALMs in long-term scenario, we construct a taxonomy of scenarios based on three essential abilities reflecting realistic long-term tool-use challenges: Context Recall, Information Shift, and Missing Context. Each scenario is further divided into two levels of difficulties (i.e.simple and complex), yielding 6 scenarios in total.
4.2.1 Context Recall (CR)
Users often refer back to previously mentioned details using pronouns or contextual ques. In this scenario, agents must ensure the continuity and relevance in dialogue by identifying and (if required) integrating relevant past information scattered across long conversations.
  • CR-Single: A single relevant piece of information (the “needle”) must be extracted from the haystack. This mirrors scenarios where a user refers to a specific past instruction or fact.
  • CR-Multi: Multiple pieces of relevant information must be identified and integrated. This setting tests the model’s ability to consolidate information from various parts of the conversation, requiring both memory and reasoning.
4.2.2 Information Shift (IS)
Long-term user interactions are dynamic—users may revise goals, introduce new constraints, or provide updated information over time. This scenario evaluates whether the agent can track these changes, even when shifts are subtle or embedded within noise. Misalignment often leads to incorrect actions, like executing outdated plans.
  • IS-Explicit: Changes are clearly stated by the user (e.g., “Actually, change the flight to May 10th instead of April 15th.”). Models must detect and adapt to these changes.
  • IS-Implicit: Shifts are implied through indirect context cues (e.g., “I won’t be able to leave that early anymore. Can you book something in the afternoon?”). This setting evaluates pragmatic understanding and contextual reasoning.
4.2.3 Missing Context (MC)
An often overlooked aspect of agent robustness is recognizing when essential information is absent (e.g., missing API parameters) and abstaining from potentially erroneous responses to avoid failures.
  • MC-Easy: Missing information is obvious, such as when no relevant entity or fact has ever been mentioned.
  • MC-Hard: The gap is subtle or masked by irrelevant but similar-seeming information, requiring fine-grained discrimination between known and unknown details.
5 Building ToolHaystack
To ensure high-quality data while maintaining cost-efficiency, ToolHaystack employs a multi-stage generation pipeline with three main phases: (1) Tool set collection; (2) Scenario-grounded tool sequence construction (3) Dialogue instantiation. Detailed process of generation process is presented in Appendix C.
Figure 3: Overview of our three-stage dataset generation pipeline.
5.1 Tool Set Collection
We begin by collecting over 13K real-world APIs from RapidAPI (RapidAPI, 2025), spanning 42 categories. Each API comes with a description of functionality and argument schema. We first sample a target API to be invoked in the target session from the curated toolset.
5.2 Scenario-Grounded Tool Sequence Construction
To avoid incoherent sessions in test instances, we construct structured tool sequences as a blueprint for subsequent generation process. This approach enables us to generate test instances with coherent long-term dependencies between tools through shared arguments.
Modeling Target–Needle Dependencies.
For a target API, we identify a set of needle APIs that can form a long-term dependency with target API through shared arguments. For example, a hotel booking API may require a location and date, provided earlier by a flight search or calendar tool.
Each argument’s dependency is identified using semantic matching (SentenceBERT (Reimers and Gurevych, 2019)) and fine-grained LLM verification to ensure the semantic relevance and compatibility. We annotate each tool call with realistic argument values using LLM prompting. Specifically, a tuple of 4 elements is given to the model i.e., target API, needle APIs, a scenario type (e.g., CR, IS, and MC), situational context. Here, situation context is introduced to enhance the authenticity and consistency of the argument values. We pre-constructed a set of situational contexts for each goal APIs based on the characteristics and requirements of each target API and randomly use one of them as a situational context.
Modeling the Haystack.
We configure a set of tools to be used in haystack (distractor) sessions based on the goal API and the selected scenario (e.g., CR, IS, and MC). An LLM is employed to compose the APIs to be used for distractors and their arguments, guided by each scenario. For instance, in MC-Easy, we use semantically disjoint tools, while in MC-Hard we include distractors that mimic needle semantics.
The prompt for second stage filtering of modeling target-needle dependencies stage and scenario grounding are presented in Appendix. We carefully craft prompts for applying each test scenario (CR, IS, MC) to build target-needle dependencies and model the haystack.
5.3 Dialogue Instantiation
With structured tool sequences fully annotated, we convert them into naturalistic dialogues. This involves rendering tool arguments into plausible user utterances and assistant responses. The needle and haystack interactions are interleaved. This results in dialogues where key information is embedded in distracting context. All instances undergo automated validation to ensure API correctness, argument consistency, and logical coherence. Then, human reviewers inspect each dialogue for naturalness and functional integrity. Only instances passing both validation stages are retained in ToolHaystack, ensuring high-quality data.
6 Experiments
6.1 Experimental Setup
Models.
We evaluate a comprehensive set of highly capable TALMs. Closed-source models include GPT family (4o, 4.5-Preview, and 4o-mini), Claude family (3.7-Sonnet, 3-Opus), Gemini Family (2.5-Pro, 2.0-Flash-001), Amazon-Nova-Pro, Mistral-large and Grok-3. Open-source models are Qwen2.5-72B-Instruct, Llama-3.3-70B-Instruct, DeepSeek-V3, ToolACE-2-Llama-3.1-8B, watt-tool-8B and xLAM-2-32b-fc-r. Note that ToolACE-2-Llama-3.1-8B, watt-tool-8B and xLAM-2-32b-fc-r are finetuned for function calling tasks. Models are selected to cover a diverse and competitive set of top-performing TALMs, particularly those demonstrating strong capabilities in recent multi-turn tool benchmarks (e.g., BFCL).
Retrieval
As providing all the API documents is infeasible, a subset of feasible APIs is typically given to TALM through semantic matching. In ToolHaystack, the model is tested in oracle setup with the 5 API documents, where a ground-truth API document and 4 APIs are given to the model based on their semantic relevance to the target session.
Method Context Recall (%) Information Shift (%) Missing Context (%) Base Avg.
Single Multi $\Delta$ Explicit Implicit $\Delta$ Easy Hard $\Delta$
Closed-source models
GPT-4o 63.44 50.00 +13.44 31.67 28.57 +3.10 21.52 20.00 +1.52 83.33 35.87
GPT-4.5-Preview 70.97 53.85 +17.12 38.33 35.71 +2.62 29.11 24.71 +4.40 72.22 42.11
GPT-4o-mini 56.99 38.46 +18.53 38.33 27.38 +10.95 29.11 12.94 +16.17 74.07 33.87
Claude-3.7-Sonnet 52.69 55.77 -3.08 25.00 21.43 +3.57 11.39 8.24 +3.15 64.81 29.09
Claude-3-Opus 34.41 30.77 +3.64 8.33 9.52 -1.19 5.06 4.71 +0.35 79.63 15.47
Gemini-2.5-Pro 48.39 57.69 -9.30 20.00 25.00 -5.00 12.66 12.94 -0.28 59.26 29.45
Amazon-Nova-Pro-v1.0 54.84 44.23 +10.61 35.00 23.81 +11.19 27.85 31.76 -3.91 52.78 36.25
Mistral-large-2407 59.14 54.08 +5.06 26.67 25.51 +1.16 20.32 23.00 -2.68 57.41 34.79
Grok-3-beta 64.52 61.54 +2.98 36.67 35.71 +0.96 15.19 20.00 -4.81 66.67 38.94
Open-source models
Qwen2.5-72B-Instruct 59.14 65.38 -6.24 31.67 28.57 +3.10 16.46 5.88 +10.58 62.96 34.52
Llama-3.3-70B-Instruct 45.16 53.85 -8.69 20.00 20.24 -0.24 0.00 0.00 +0.00 74.07 23.21
DeepSeek-V3 55.91 59.62 -3.71 40.00 28.57 +11.43 27.85 18.82 +9.03 68.52 38.46
ToolACE-2-Llama-3.1-8B 23.66 17.31 +6.35 10.00 4.67 +5.33 0.00 0.00 +0.00 70.37 9.27
watt-tool-8B 15.05 5.77 +9.28 8.33 2.38 +5.95 0.00 0.00 +0.00 62.96 5.26
xLAM-2-32b-fc-r 50.62 51.06 -0.44 26.92 25.33 +1.59 0.00 0.00 +0.00 69.81 25.66
Table 2: Main results. $\Delta$ denotes the gap between simple and complex levels. Base denotes BFCL score; Avg. is the mean performance across metrics.
Evaluation Metric.
We use call accuracy which calculates the proportion of correct function calls:
Call $\text{Call Acc}=\frac{\sum_{i=1}^{N}c_{i}}{\sum_{i=1}^{N}n_{i}}$ (1)
where $N$ is the number of samples, $n_{i}$ is the total number of function calls in $i$-th sample, and $c_{i}$ is the number of correct function calls in sample $i$.
Base Test Set (Multi-turn).
To provide deeper insight, we generate a multi-turn counterpart of our dataset generated by standard self-instruct approach previously used in the literature (Qin et al., 2023). We use the same goal APIs for generating short multi-turn tool-use instances. All necessary information to complete the task is revealed within a few turns, without any distractions.
6.2 Main Results
We present the ToolHaystack results in Table 2, evaluating 17 TALMs across six long-term interaction scenarios.
Overall Performance.
Among all models, GPT-4.5-Preview achieved the highest average score (42.11%), followed by Grok-3-beta (38.94%) and DeepSeek-V3 (38.46%). While current TALMs show competitive performance in multi-turn settings (i.e., Base), their accuracy significantly deteriorates in ToolHaystack. For instance, GPT-4o drops from 83.33% in Base to 35.87%, demonstrating the complexity of long-term tool-use.
Noise Robustness.
The performance gap between simple and complex sub-scenarios shows the robustness of models against contextual noises. We observe GPT family models struggles to retain their performance in CR-Single and CR-Multi. For example, GPT-4.5-Preview, one of the strongest model in ToolHaystack, underperform significantly in CR-Multi. In contrast, Claude family models and Grok-3-beta have experienced marginal performance drops by contextual noises, showing under 5% of degradation across scenarios consistently.
Closed- vs. Open-source Models.
Closed-source models consistently outperform open-source models under noisy settings. While general-purpose open-source models (Qwen2.5, DeepSeek-V3) demonstrate more stable performance across scenarios, they still lag behind top proprietary models. Notably, open-source finetuned models like ToolACE and watt-tool achieve generally achieve poor accruracy of less than 10%, despite their high BFCL scores. We further examine this by extending the comparison between multi-turn and long-term tool-use performance using BFCL benchmark in the following section.
6.3 Limitation of Multi-turn Benchmarks for Evaluating Long-term Tool Use
To investigate whether existing multi-turn benchmarks can reliably reflect long-term tool use performance, we compare performance on BFCL (multi-turn) and ToolHaystack (long-term). As shown in Figure 4, a positive correlation between the two benchmarks is observed within each group (general-purpose LLMs and finetuned models). However, while considering all models, we find a notable discrepancy: finetuned models achieve high scores on BFCL but perform poorly on ToolHaystack. This gap suggests that multi-turn benchmarks may fail to capture the complexity and robustness required for long-term interactions. One possible explanation is an out-of-distribution (OOD) problem—models trained on conventional tool-learning corpus struggle to generalize to diverse, noisy and unpredictable nature of real-world scenarios. As fine-tuning TALM is considered a promising direction to build personal agents, the lack of generalization in long-term tool-use can be a critical issue, highlighting the importance of evaluating tool use capabilities of LLMs during long-term interaction such as ToolHaystack.
Figure 4: Performance comparison between BFCL (multi-turn) and ToolHaystack (long-term) score.
Distance 0 (recent) 1 2 3 4 5 (past) Closed 94.44 41.67 61.11 75.00 44.44 55.56 General 83.33 25.00 50.00 66.67 33.33 58.33 TALM 33.33 0.00 33.33 41.67 25.00 50.00
Table 3: Accuracy by evidence position (distance from evidence) across closed-source models, general-purpose open-sourced models, and finetuned TALMs. The numbers are averaged scores of each model group.
6.4 Controlled Analysis on Long-Context Utilization of TALMs
To isolate the impact of long-context challenges in tool-augmented settings, we conduct a controlled ablation study by systematically modifying the contextual structure in ToolHaystack.
Positional Bias in TALMs.
We analyze how model performance varies with the position of the key evidence ("needle") in the context. We report the averaged accuracy of models of three groups of models, namely Closed-source models, Open-source general purpose models, and finetuned TALMs, as a function of the number of sessions between the session containing the relevant evidence and the target session. As shown in Table 3, closed-source models exhibit a strong recency bias, achieving peak performance (94.44%) when the evidence is immediately adjacent to the query (distance 0). General-purpose open-source models follow a similar pattern but with lower overall robustness. In contrast, finetuned TALMs struggle across all positions. Overall, the findings underscore that positional biases significantly affect tool invocation accuracy.
Figure 5: Performance degradation occurs as the number of distractor haystacks increases.
Impact of Distractors.
We investigate how increasing the number of haystack sessions lead to decreased performance. We select top-performing TALMs in ToolHaystack and assess them on the samples with varying number of haystacks. In Figure 5, model performance consistently declines as the number of haystacks increases. For example, GPT-4.5-Preview’s accuracy drops sharply from 56% to 32% when the number of distractors grows from 1 to 20. Interestingly, the performance gap between GPT-4.5 and other models narrows at higher noise levels, suggesting a saturation point beyond which all models are overwhelmed by irrelevant context.
6.5 Error Analysis through the Lens of Hallucination
As interaction history accumulates over time, TALMs are increasingly prone to hallucinating arguments or selecting incorrect tools by referring to irrelevant yet semantically-related context. We investigate model failures in the Missing Context (MC) scenarios where current TALMs struggle the most by analyzing the distribution of error types across model classes. Figure 6 presents the relative proportions of three key error types: wrong API selection, hallucinated arguments based on misleading in-context cues (In-context), and hallucinations derived from entirely absent information (Out-of-context). The numbers are averaged within each model group.
Figure 6: Compact error breakdown by model and context shown as percentages (API: wrong API name, In/Out-of-context: wrong parameter values filled with entities appearing in/out of context).
Error Trends.
Closed and General-purpose models exhibit highly skewed error distributions, with over 75% of their failures attributable to Out-of-context hallucinations in MC-Easy and similarly high proportions in MC-Hard (63% and 77%). We find that several models tend to copy default values of missing parameter or filling out randomly generated strings without proper verification of contextual sufficiency. Finetuned TALMs, in contrast, show a different error distribution. In MC-Easy, TALMs reduce Out-of-context hallucinations to 43.39% and In-context hallucinations remain low (7.94%), while API errors account for nearly half (48.68%) of the total errors, suggesting that TALMs are more cautious about fabricating arguments outright from missing context (lower Out-of-context) but often misidentify the appropriate tool (API error), particularly under harder conditions. See Appendix D.1 for detailed case studies illustrating model behaviors.
6.6 Does CoT Prompting Elicit Better Performance?
We investigate whether CoT also improves robustness in long-term tool-use settings, where contextual entanglement, evolving goals, and noisy discourse severely challenge model reliability.
Setup.
We conduct a controlled comparison of five models showing competitive performance in our dataset—GPT-4o, GPT-4.5-Preview, Amazon-Nova-Pro-v1.0, Grok-3-beta, and DeepSeek-V3 across the test scenarios in ToolHaystack.
Method CR (%) IS (%) MC (%) Avg.
S M E I E H
GPT-4o 63.44 50.00 31.67 28.57 21.52 20.00 35.87
+ CoT 58.06 69.23 30.00 30.95 18.99 16.47 37.95
GPT-4.5-Preview 70.97 53.85 38.33 35.71 29.11 24.71 42.11
+ CoT 70.97 61.54 30.00 42.86 31.65 22.35 43.23
Amazon-Nova 54.84 44.23 35.00 23.81 27.85 31.76 36.25
+ CoT 52.69 51.92 25.00 26.19 17.72 25.88 33.23
Grok-3-beta 64.52 61.54 36.67 35.71 15.19 20.00 38.94
+ CoT 59.14 71.15 38.33 36.90 13.92 23.53 40.83
DeepSeek-V3 55.91 59.62 40.00 28.57 27.85 18.82 38.46
+ CoT 51.61 71.15 36.67 33.33 17.72 15.29 37.63
Table 4: The results of each method are evaluated using the same setup as the main results. Rows listing only the model name correspond to the vanilla (zero-shot) setting. The detailed prompts for CoT are in the Appendix. Note that Amazon-Nova stands for Amazon-Nova-pro-v1.
Results.
Our findings, summarized in Table 4, reveal mixed effects of CoT prompting. While CoT often yields marginal or scenario-specific improvements, it does not consistently improve overall performance (e.g.GPT-4o). Among the models, GPT-4.5-Preview and Grok-3-beta shows the most robust gains, particularly in CR-Multi and IS-Implicit, suggesting its stronger alignment with reasoning-based cues. DeepSeek-V3, however, sees limited benefit, possibly due to the quality of generated reasoning paths. These underscore that CoT prompting is not a universally effective for long-term tool use. Improvements are scenario-dependent and closely tied to the model’s inherent reasoning capabilities and contextual retrieval mechanisms.
7 Conclusion
In this work, we introduced ToolHaystack, a novel benchmark designed to rigorously assess the long-term robustness of tool-augmented language models (TALMs) in complex, noisy, and realistic interaction scenarios. Our extensive evaluations across 10 state-of-the-art models demonstrate that while existing TALMs excel in short, clean multi-turn dialogues, their performance significantly degrades in long-term, distraction-rich contexts. This performance gap highlights the critical need for dedicated long-term evaluation frameworks and reveals current limitations in models’ ability to retain context, handle goal shifts, and avoid hallucination over time. ToolHaystack offers a composable, scalable, and diagnostically rich test suite to drive progress toward more reliable, agentic LLM systems. We hope our benchmark facilitates future research aimed at closing the gap between controlled evaluations and the complex realities of real-world tool use.
8 Limitations
Analysis on the Reasoning-Intensive Models.
Recently, large reasoning models (LRMs), which are trained to improve reasoning through the use of extended rationales, have demonstrated strong performance across a range of tasks. However, it remains unclear whether such reasoning capabilities directly translate to effective tool use in long-term interactions, which often require pragmatic reasoning over extended contexts. While LRMs may excel at structured reasoning within a single context window, adapting to evolving goals and maintaining coherence across multiple tool-use steps poses additional challenges. Investigating this connection is a promising direction for future work.
Evaluation on Extremely Long-term Interactions.
In real-world applications, TALMs may interact with users over extended periods, ranging from months to years. To maintain evaluation quality and manageability, our current setup limits the number of sessions. Nevertheless, a valuable future direction is to scale our datasets to hundreds of sessions, enabling evaluation of in-the-wild usage patterns and better assessing TALMs’ ability to maintain consistency, memory, and utility over long-term interactions.
Number of Used APIs.
Although our dataset includes a limited number of APIs, our aim is to focus on those that are properly documented, functional, and realistic for use in real-world applications. During the exploration phase, we observed that incorporating APIs that did not meet these criteria significantly degraded data quality. While one could manually construct additional high-quality APIs to expand ToolHaystack, such efforts are beyond the scope of this work.
Appendix
Appendix A Detailed Related Work
A.1 Tool-Augmented Language Models
Large Language Models (LLMs) augmented with external tools have emerged as a promising paradigm for overcoming inherent limitations in static parametric memory (Komeili et al., 2022) and enabling real-time interaction with the external world (Yang et al., 2024; Chae et al., 2025). The term tool encompasses a broad range of functionalities, from simple calculators and Python interpreters (Cai et al., 2024) to complex API and function calls (Liu et al., 2024). In this work, we focus specifically on the latter—LLMs interfacing with external APIs or invoking function calls to perform tasks.
Early efforts concentrated on enabling LLMs to call tools via prompting and reasoning frameworks. For instance, Parisi et al. (2022) introduce the concept of Tool Augmented Language Models (TALM), demonstrating that an LLM can iteratively interact with non-differentiable tools through textual interfaces. Similarly, the ReAct framework (Yao et al., 2022) combined chain-of-thought reasoning with action directives, allowing an LLM to plan steps and invoke tools (e.g., web search or calculators) in a conversational loop. Subsequent research has shown that fine-tuning LLMs on tool-use data can significantly enhance their efficacy. Toolformer (Schick et al., 2023), for example, is a fine-tuned model with self-generated tool API calls inserted into its training data, enabling it to decide when and how to use calculators, search engines, translators, and other APIs.
With the rising demand for practical task-solving capabilities, recent proprietary LLMs have evolved to support tool or function calling as a core interface. Starting with GPT-4’s function calling API (OpenAI, 2023), OpenAI has progressively enhanced tool interoperability in their models, culminating in GPT-4.1 and GPT-4o, which natively support multimodal input, extended context windows, and robust tool invocation within conversation flows. Similarly, models like Claude 3.5 (Anthropic, 2024) and Gemini 1.5 (Team et al., 2024) offer built-in support for external API access and tool orchestration, enabling seamless integration of search, calculator, code execution, and retrieval systems in their agent pipelines. These developments mark a paradigm shift from pure text generation to real-world interaction, where tool use becomes a first-class citizen in the model’s interface. Alongside model-level enhancements, ecosystem-level tool agents have also proliferated. For example, OpenAI’s ChatGPT integrates plugin systems for third-party APIs and file browsing, while Google’s Bard/Gemini and Anthropic’s Claude provide tool-use scaffolding through memory and workflow APIs. These systems often combine planning, execution, and memory modules, highlighting the emerging need for robust, API-centric LLM benchmarks and evaluation methods.
Recognizing the performance gap between open-source LLMs and proprietary models in tool calling, recent efforts have focused on equipping smaller or open LLMs with competitive tool-use abilities. For instance, Gorilla (Patil et al., 2023) fine-tunes LLaMA models on a large corpus of API documentation and call traces, leveraging a retriever-generator architecture to improve accuracy and reduce hallucination. ToolLLM (Qin et al., 2023) adopts a reasoning-guided planning strategy that explores multiple API invocation paths via a depth-first search over possible tool sequences. Salesforce’s xLAM (Zhang et al., 2024) provides a family of open LLMs specifically tuned for tool usage, achieving strong performance on multi-tool benchmarks. Other open initiatives such as ToolAlpaca (Tang et al., 2023), AnyTool (Du et al., 2024), and ToolBridge (Jin et al., 2024) introduce simulated training pipelines or meta-learning objectives for improving tool generalization in smaller models. These efforts aim to bridge the gap in real-world utility and provide a competitive open-source alternative to commercial tool-augmented LLM systems.
A.2 Benchmarks for Evaluating TALMs
As TALMs have advanced, various benchmarks have been introduced to evaluate their capabilities in using external APIs. Early benchmarks primarily assessed models in static, single-turn settings without environmental feedback. For instance, API-Bank (Li et al., 2023) provides 314 task scenarios across 73 APIs, measuring planning, selection, and calling accuracy. Similarly, APIBench (Patil et al., 2023) evaluates whether a model can generate syntactically and semantically correct API calls over a large corpus of public functions. These benchmarks test the model’s symbolic reasoning and API schema understanding but lack interactive components.
Subsequent work introduces real-world interaction into the evaluation loop, enabling models to be assessed in dynamic environments. RestGPT (Song et al., 2023) connects LLM agents to actual RESTful APIs and evaluates their ability to plan API sequences and recover from execution errors. To further approximate user-facing scenarios, recent benchmarks have introduced user simulators and multi-turn dialogues. ToolDial (Shim et al., 2025) includes over 11k dialogues annotated with 16 action types, capturing realistic tool-assisted assistant behavior with clarifications, failures, and API chaining. $\tau$-bench (Yao et al., 2025) extends this idea by simulating domain-constrained interactions (e.g., shopping, airline booking) and measuring goal completion in a transactional database, along with a consistency metric (pass@$k$) to evaluate reliability across repeated trials.
Recent benchmarks have introduced more complex setups to test advanced capabilities like planning and tool use. UltraTool (Huang et al., 2024) separates planning from execution to assess models’ ability to devise tool-use plans. ToolSandbox (Lu et al., 2024) adds stateful tools and on-policy user simulators, supporting persistent states and multi-tool chaining. HammerBench (Wang et al., 2024) expands on this with diverse multi-turn tasks, including Q&A and argument shifts. ComplexFuncBench (Zhong et al., 2025) addresses the challenge of extremely long contexts (e.g., 128k tokens) and proposes constructing long function call sequences.
Despite these developments, a key missing piece in existing benchmarks is support for multi-session evaluation—tracking user preferences, goals, and tool interactions across sessions that reflect the long-term interaction between users and TALMs. Our work addresses this gap by introducing a benchmark that explicitly evaluates TALMs in multi-session settings, enabling more realistic measurement of memory, personalization, and tool-use generalization over time.
Appendix B Dataset Details
Rank Category # of APIs Percentage
1 Data 1522 11.48%
2 Social 926 6.98%
3 eCommerce 845 6.37%
4 Other 838 6.32%
5 Finance 732 5.52%
6 Business 715 5.39%
7 Entertainment 690 5.20%
8 Sports 677 5.11%
9 Music 514 3.88%
10 Business_Software 423 3.19%
11 Gaming 397 2.99%
12 Location 369 2.78%
13 Education 361 2.72%
14 Travel 361 2.72%
15 Email 356 2.68%
16 Communication 301 2.27%
17 Database 299 2.25%
18 Financial 290 2.19%
19 Tools 287 2.16%
20 Commerce 287 2.16%
21 News_Media 246 1.86%
22 Weather 236 1.78%
23 Food 229 1.73%
24 Media 218 1.64%
25 Movies 141 1.06%
26 Search 125 0.94%
27 Health_and_Fitness 124 0.94%
28 Science 113 0.85%
29 Text_Analysis 95 0.72%
30 SMS 92 0.69%
Table 5: Top-30 API categories with number of APIs and percentages.
B.1 Tool Set
To construct realistic tool-calling conversations, we collect real-world APIs from RapidAPI from various domains such as travel, finance, and messaging. We provide the statistics of the APIs used for our dataset construction in Table 5. As shown in the table, we utilize various APIs across multiple domains.
Metric Context Recall Information Shift Missing Context
Single Multi Explicit Implicit Easy Hard
# Examples 93 52 60 84 79 85
Utterance Tokens Mean 1440.71 1450.15 1554.83 1542.93 878.70 1339.56
API Tokens Mean 6913.02 8677.83 7798.22 10797.85 6372.62 9014.56
Total Tokens Mean 8353.73 10127.98 9353.05 12340.77 7251.32 10354.13
Sessions Mean 16.08 13.33 14.22 13.58 11.30 12.76
Turns Mean 76.77 69.94 76.15 78.80 50.18 64.52
Table 6: Statistics of ToolHaystack
B.2 Dataset Statistics
The statistical overview of ToolHaystack is provided in Table 6. The dataset comprises three main categories, each split into easy and hard settings. For every category, we include over 100 examples that have been strictly vetted through our filtering pipeline. As demonstrated in Figure 7, ToolHaystack features a wide range of token distributions, allowing for robust testing of long-term interaction performance in existing LLMs.
Figure 7: Token length distribution of ToolHaystack
Figure 8: Real-world interactions between human and agent are intertwined, with natural contextual noise accumulated over time.
Appendix C Details in ToolHaystack Construction
C.1 Tool Collection
We collect over 13K real-world APIs from RapidAPI [^3] https://rapidapi.com/, across domains such as travel, finance, and messaging. APIs lacking stable endpoints or adequate documentation are filtered out. For each selected API, we extract method descriptions, parameter lists, and usage examples. These are normalized into a consistent format using GPT-4o-mini to support downstream tool planning and prompt generation.
API Name Scenario Description
Search Plant By ID A landscape architect from Brazil moves to Canada for a project focusing on sustainable urban gardens. While researching native flora, she needs to identify a specific plant by its ID to ensure compatibility with the local ecosystem.
An aspiring botanist in the UK volunteers at a community garden, where he discovers a rare plant species. Curious about its properties and care requirements, he looks it up using the API to provide accurate information to the garden’s committee.
A retired school teacher in Australia starts a blog on native plants after moving from the USA. She wants to showcase a particular plant she used to teach about, so she searches for its ID through the API to gather detailed information.
Table 7: Example scenarios for the Search Plant By ID API
C.2 Modeling Tool Sequences
At the core of each instance lies a tool sequence—a directed list of API calls encoding long-term dependencies. The final target call reuses arguments or outputs from prior needle calls. For example, a hotel booking API might require location and check-in date, both previously produced by a flight search or calendar lookup.
We begin by selecting a target API and then identify prior APIs whose outputs semantically align with its required arguments. This alignment is computed using embedding-based similarity (e.g., Sentence-BERT), followed by LLM-based filtering to remove misleading matches such as unrelated IDs. For each target argument, we annotate whether its source is a previous tool call, user utterance, or remains unresolved (used later in missing context scenarios).
Each sequence is grounded in a plausible situation (e.g., vacation planning, delivery tracking) generated by prompting LLMs to create natural task flows. This provides contextual motivation for tool usage and ensures that dependencies are meaningful and realistic.
For situation generation, we prompt gpt-4o-mini model to generate 30 plausible scenarios related to the API information to gather diverse situation and for each sequence, we randomly sample a single situation from 30 scenarios. We provide generated situation examples for a random API in Table 7.
C.3 Scenario-Grounded Tool Sequence Construction
To introduce specific interaction challenges, we embed each tool sequence into one of the ToolHaystack scenarios. For instance, CR scenarios scatter key arguments across earlier dialogue turns; IS scenarios modify goals mid-way through the interaction; MC scenarios mask required arguments to test the model’s ability to ask for clarification. These transformations are applied by prompting LLMs to rewrite the original dialogue into one that matches the structural requirements of the scenario, while preserving functional correctness of the tool calls.
Based on the scenario, we first select the proper APIs that will be included in the tool sequence. For example, for ContextRecall scenario, we choose the API pair with similar parameter(s) as input from the previously aligned API set. Here, for ContextRecall-Multi scenario, we make the API triplet where two APIs share a single argument with the goal API.
C.4 Modeling Haystack
To simulate long-term, noisy conversations, we interleave the tool-relevant dialogue with distractor sessions—unrelated sub-dialogues involving different user intents or casual exchanges. These distractors are independently generated and inserted before, between, or after the tool-relevant content. The number, type, and semantic distance of these distractors are controlled to tune the difficulty of the instance. The result is a single, contiguous dialogue in which critical information is sparsely and irregularly distributed, mimicking the cluttered nature of real-world interactions.
For modeling haystack, we provide previous generated tool sequence as input and randomly select the APIs that can be utilized as haystack. Here, as haystack should not affect the evaluation, we filter out APIs by checking if the randomly selected APIs have any argument that has similar functionality with any argument in the APIs in tool sequence.
C.5 Interaction Simulation
Agent-User Interaction
Using the tool sequence and scenario specification, we simulate a complete user–agent conversation. The user progressively reveals goals and arguments, while the agent responds through tool calls or natural replies. Each turn is aligned with a structured plan, preserving traceability between user inputs, system actions, and tool responses.
C.6 Quality Assurance
We employ multiple frontier LLMs (GPT-4o, Claude 3.5, Gemini 2.5, LLaMA 3.3) during generation to reduce model-specific artifacts. Automated validation enforces constraints such as correct argument use, proper tool syntax, and scenario fidelity. A subset of generated dialogues undergoes a filtering process, where we prompt gpt-4o to strictly evaluate if the dialogue violates the rules for the target category.
Moreover, to ensure the quality of ToolHaystack, we employed GPT-4o as a judge to automatically filter out dialogues that do not satisfy our scenario definitions. To validate the effectiveness of this automated filtering process, we sampled 60 dialogues (10 from each category) from the unfiltered dataset and compared the results against human annotations. The evaluation yielded a Matthews Correlation Coefficient (MCC) of 0.884, indicating a strong agreement between the automated and human filtering outcomes.
C.7 Dataset Generation Details
For ContextRecall-$multi$ scenario, we find that if APIs used for needle include multiple parameters that work similarly with any parameter in the target API, it can repeatedly appear in the dialogue context and makes it easier to find the needle. To this reason, we utilize API triplets where APIs used for needle have a single shared argument with the target API.
For InformationShift scenario, we find that if multiple arguments shifts over turns, most of models fails to generate the correct API call. To this reason, in this scenario, only a single argument evolves in multi-turn conversations.
For MissingContext-$Hard$ scenario, the user query mentions the argument that is needed in detail to prevent the model being confused with other distractor arguments. For example, if the user asked to use OpenAI API in previous session with API key and asks to use Anthropic API without mentioning API key, the model can be confused to use openAI API key. For this situation, we prompt the model to generate detailed query like “Please use the same API key that I used for Anthropic API previously” instead of asking “Please use the same API key that I used”.
Appendix D Evaluation Details
We use greedy decoding (i.e., temperature set to 0) and limit the maximum number of generated tokens to 2,000. For closed-source models, we access the GPT family via the OpenAI API and the Claude models via the Anthropic API. Other models accessible through OpenRouter are used via that service. Finetuned TALMs are served using vLLM (Kwon et al., 2023).
Each model requires specific input formatting; for example, some expect tool descriptions to be provided in JSON rather than Markdown. We follow the formatting guidelines provided in each model’s official documentation. All experiments are conducted using three NVIDIA A100 80GB GPUs. All evaluations are conducted in a single run.
D.1 Case Study
Table 8 shows an example failure of GPT-4o-mini.
Sample Dialogue from Missing Context-Hard:
(Previous sessions)
User: Hey, I need to check the domains hosted on the same IP as a suspicious domain. Use the domain I mentioned before, and let’s go with the page number we talked about last time.
Ground-truth Response:
  "api_name": "Search domains",
  "parameters": {
    "q": "suspiciousdomain.com",
    "type": "WEBIP",
    "pagenum": "MISSING"
GPT-4o-mini Prediction:
  "api_name": "Search domains",
  "parameters": {
    "q": "suspiciousdomain.com",
    "type": "WEBIP",
    "pagenum": "3"
DeepSeek-V3 Output:
  "api_name": "Search domains",
  "parameters": {
    "q": "suspiciousdomain.com",
    "type": "WEBIP",
    "pagenum": "1"
Table 8: An example from the Missing Context-Hard setting showing the dialogue, ground-truth response, and model predictions. GPT-4-o-mini predicted pagenum as "3" which is a random number and DeepSeek-V3 predicted pagenum as "1" which is a default value.
Prompt for Dialogue Generation (CR-S)
Prompt for Dialogue Generation (CR-M)
Prompt for Dialogue Generation (IS-E)
Prompt for Dialogue Generation (IS-I)
Prompt for Dialogue Generation (MC-E)
Prompt for Dialogue Generation (MC-H)
Prompt for Filtering Dialogue
Prompt for Filtering Tool Sequence (CR-S)
Prompt for Filtering Tool Sequence (CR-M)
Prompt for Filtering Tool Sequence (IS)
Prompt for Filtering Tool Sequence (MC)
Prompt for Generting Distractor
Prompt for Evaluation